What equation can I use for the following: Melissa has 5 times as much money as Manuel. They have $60.00 together?
Answer by sdmmadam@yahoo.com(530)

Tutoring Home For Students Tools for Tutors Our Tutors Register to be a tutor Recently Solved


Question 29097 in Evaluation_Word_Problems:
Melissa has 5 times as much money as Manuel. They have $60.00 together. What equation can I write for this?



Melissa has 5 times as much money as Manuel. They have $60.00 together. What equation can I write for this?
Let Melissa have L$
and let Manuel have M$
Melissa has 5 times as much money as Manuel.
That is L= 5 M----(1)
They have $60.00 together
That is L+M= 60 ----(2)
Use (1) in (2)
5M+m=60
6M= 60
M=60/6= 10
M=10 in (1) gives L=5M = 5X10= 50
Answer: Melissa has 50 dollars and Manuel has 10 dollars
Verification: 50 is of course 5 times 10 and 50 and 10 are together 60